# Task-Augmented Cross-View Imputation Network for Partial Multi-View Incomplete Multi-Label Classification

### v1

@@ Abstract
task-augmented, cross-view imputation, partial multi-view learning, incomplete multi-label classification.
I Introduction
@@
In recent years, with the significant progress of representation learning methods, traditional single-view data analysis methods are gradually showing their limitations, and they are difficult to fully adapt to the growing demands of diverse application scenarios [1]. Therefore, multi-view data integration has become an important way to break through the inherent barriers of single data sources and alleviate the lack of information. This trend is accompanied by the rapid development of data acquisition technology, which makes a huge amount of multi-view data from all kinds of media or with different styles emerge, and providing more possibilities for achieving a comprehensive and accurate description of observation targets [2][3]. In short, multi-view data refers to a series of complementary information collections collected from different observation angles for the same object [4][5]. For example, three unique visual information obtained from X-rays, CT scans, and magnetic resonance imaging (MRI) in medical diagnostics constitute a multi-view dataset; retinal images acquired at four different positions can also constitute another multi-view retinal dataset [6][7]. More commonly, multi-view or multi-modal datasets that incorporate multiple types of data including text, images, and videos have been utilized in many application scenarios such as web page classification [8][9]. Concurrently, it is observed that single-label data frequently fail to adequately reflect the intricacies of real-world scenarios [10]. For instance, a picture of a flower may also require labels such as “winter” or “greenhouse”, demonstrating that multi-label provide a more comprehensive category feature space and more accurately preserve inter-label relationships. Given the diversity of objects in multi-label classification task, researchers have innovatively combined multi-view learning with multi-label classification, making multi-view multi-label classification (MVMLC) a domain of substantial research interest [11].
@@
For MVMLC, many methods have been proposed in the past few years, such as the latent semantic aware MvMLC [12], multi-view probabilistic model LCBM [13], and maximizing label-feature dependency based method [14]. However, MVMLC assumes all views and their corresponding labels are complete, a premise that does not always correspond to real-world situations. For instance, in the context of video material, audio tracks are often absent, and text descriptions may be incomplete. Similarly, multimedia content disseminated on social media platforms frequently exhibits incomplete annotations owing to varying levels of user engagement. These practical situations give rise to the common challenge of incomplete multi-view features and incomplete category information in real-world datasets, which we define in our paper as “partial multi-view incomplete multi-label classification (PMVIMLC)” [15][16]. To solve such problem, various approaches for PMVIMLC have been proposed recent years. Tan et al. initially introduced the iMvWL model to construct a shared low-dimensional subspace based on matrix decomposition [11]. And Li et al. proposed the NAIM3L model to preserve the global consistency and local properties of the label structure [17]. Nevertheless, since these models employ shallow machine learning techniques to explore the information contained in features and labels, they are constrained in their ability to extract higher-order discriminative features from data [9]. Therefore, Wen et al. employed deep neural networks to design the DD-IMvMLC-net model while Liu et al. developed the DICNet model for extracting deeper semantic representations from multi-view data [9][18]. However, the aforementioned models introduce a missing view indication matrix to circumvent unknown views and rely solely on the available data for classification, which fails to fully utilize the potential cross-view complementary information and the intrinsic relationships of the missing data, rendering it suboptimal [19]. To address this limitation, we propose to construct models that are capable of directly recovering missing data, despite the inherent challenges associated with this task.
@@
In light of such an issue, we propose a task-augmented cross-view imputation network (TACVI-Net) for partial multi-view incomplete multi-label classification. TACVI-Net has been designed to handle arbitrary view-missing scenarios and enhance representation learning through effective view completion. It is acknowledged that raw data may contain task-irrelevant redundant information, which may result in the accumulation of such information if used directly to fill in missing data. Therefore, we design a view-specific encoder-classifier architecture based on variational coding, inspired by the information bottleneck theory. This architecture can obtain enhanced features that are highly relevant to the task so that effectively avoiding the impact of irrelevant redundant information on model performance. In the absence of view features, it is of the utmost importance to determine an appropriate filling strategy to provide sufficient and useful information. Given the existence of the shared semantic information across multi-view, we convert each input view into a hidden representation via the view encoder and merge them into a joint representation to form a fused shared representation. Subsequently, the fused shared representation is employed to reconstruct missing data via the view decoder, which is then integrated with the existing data to construct a complete dataset. In this way, we can obtain more sufficient information to exploit the consistency and complementarity across different views, thus improving the final classification performance.
@@
TABLE I: Main notations utilized in our paper.
@@
$Y^{{}^{\prime}}$ predictive label matrix ($(\left\{0,1\right\}^{n\times c})$
@@
$\Theta$ classifier for the second stage of fusion feature $\widehat{Z_{T}}$
II Preliminaries
II-A Problem Statement and Associated Symbols
@@
For the PMVIMLC task, we delineate the following notations: $\left\{X_{l}\right\}_{l=1}^{m}$ and $Y$ symbolize the original multi-view data and corresponding labels, respectively, with $m$ denoting the total number of views. Each $X_{l}=\left\{x_{l_{1}},x_{l_{2}},...,x_{l_{n}}\right\}\in R^{n\times d_{l_{x}}}$ encapsulates the properties of instances within the $l$-th view, where $n$ represents the instance count and $d_{l_{x}}$ specifies the feature dimensionality for this view. Specifically, $x_{l_{i}}$ denotes the attributes of the $i$-th sample observed through the $l$-th perspective. The label matrix $Y=\left\{y_{1},y_{2},...,y_{n}\right\}\in\left\{0,1\right\}^{n\times c}$, with $c$ represents the label count, and $Y_{i,j}=1$ implies the $i$-th instance carries the $j$-th tag; conversely, $Y_{i,j}=0$. To further account for missing data, we introduce a missing-view indication matrix $U\in\left\{0,1\right\}^{n\times m}$, where $U_{i,j}=1$ affirms the presence of the $i$-th sample in the $j$-th view, otherwise $U_{i,j}=0$. Similarly, $G\in\left\{0,1\right\}^{n\times c}$ serves as the missing-label indication matrix, with $G_{i,j}=1$ signifying awareness of the $j$-th label for the $i$-th instance, or $G_{i,j}=0$ for unknown status. In preprocessing, missing values in feature matrices $\left\{X_{l}\right\}_{l=1}^{m}$ and label matrix $Y$ are set as 0. Given the above complete or incomplete multi-view training data with complete or incomplete labels, the objective of our PMVIMLC task is to train a model that can accurately predict complete label sets for each multi-view sample with complete or incomplete views. To facilitate readers’ rapid identification of essential notations, we present a comprehensive list of the primary notations used in this paper in Table I.
II-B Related Works
II-B1 Multi-view multi-label classification
@@
The field of MVMLC focuses on coping with data that contain multiple views and multiple labels at the same time which significantly increasing the complexity. Incorporating multi-view learning strategies has been shown to be more effective in solving such problems compared to multi-label classification approaches that rely only on a single view, inspiring a large increase of related research. For instance, Zhang et al. proposed a multi-view multi-label learning method based on matrix factorization to exploit the complementarities between different views [12]. Wu et al. introduced a method called SIMM, which employs orthogonal constraints on the shared subspace to extract unique information from multiple views, simultaneously optimizing the adversarial loss and the multi-label loss to obtain shared information across views [20]. Zhang et al. introduced RAIN, which facilitates interaction and semantic label embedding under the mechanism of multiple heads of attention to enhance label associations and dynamically acquire specific label representations based on learned label embedding [21]. Sun et al. put forward the generative model LCBM to fuse multi-view features and establish inter-label dependencies by using a Bernoulli mixture model [13].
@@
The CDMM method proposed by Zhao et al. employs a nonlinear kernel neural network to create separate classifiers for each view, ensuring inter-view consistency, and incorporates the Hilbert-Schmidt criterion to enhance inter-view diversity [22]. Yu et al. presented the M3AL method, which combines multi-view self-representation learning to optimize information separation, and employs a unique query strategy to select highly informative sample-label pairs for active learning based on shared, individual information and view differences [23]. Zhao et al. introduced the LVSL method for non-aligned multi-view multi-label classification, aiming to capture view-specific label information and low-level label structure under a unified system. This method exploits structural consistency between views and labels and enhances label spatial information [24]. The D-VSM approach proposed by Lyu et al. employs deep GCN, eliminating the need for shared subspace identification. Instead, feature representations are directly encoded at the individual view level using a deep graph convolutional network, facilitating the natural coupling of information between views [25]. Yin et al. put forward the MMDOM method, which ensures that the low-dimensional shared subspace maintains both local and global data structures through first-order and second-order similarity matrices, effectively protecting manifold structures [26]. Xiao et al. developed an SVM-based approach, MVMLP, to obtain consensus information by reinforcing output similarity between views while introducing complementarity by using different views as privileged information for each other [27].
II-B2 Partial multi-view incomplete multi-label classification
@@
Wen et al. introduced DD-IMvMLC-net, a neural network-based model designed for extracting high-level features. It consolidates multi-view data through a weighted fusion approach to construct a comprehensive sample feature representation, facilitating classification [18]. Building upon this, Liu et al. advanced DICNet, which incorporates contrastive learning into the multi-view learning framework, enhancing the model’s capacity for information extraction across diverse views [9]. Liu et al. also set forth the MTD model, which decouples the common single-channel view-level representation in multi-view deep learning methods into shared and view-specific representations [31]. Tan et al. advocated for the application of information theory as a means to enhance and refine datasets that are inherently incomplete. The theoretical framework adopted by the researchers allows them to balance the available information to emphasize what is crucial for the task [32]. The SIP model, proposed by Liu and Xu et al., attempts to compress cross-view representations to maximize the amount of shared information, thereby better predicting semantic labels [33]. The above methods of avoiding or eliminating missing data are effective but not a fundamental solution. Therefore, we aim to go beyond avoidance strategies and explore new method to recover missing data.
Figure 1: The sketch of our TACVI-Net. In our network, the task-augmented encoder (TAEncoder) in the first stage is a view-specific encoder founded on variational coding. This encoder is trained in conjunction with the view-specific classifier, aimed at extracting features that optimally align with the classification objective. The second stage employs an MLP-based view-specific encoder (VSEncoder) and view-specific decoder (VSDncoder). This stage fills in the missing view features, and then inputs the supplemental multi-view data into the classification module, thus realizing the process from feature extraction to information integration, and finally to an efficient accurate classification decision-making process.
III The proposed method
@@
III-A Motivation
For one thing, it is well known that different views share the same high-level semantic information in the classification task because they are different descriptions of the same abstract target. This means that it becomes feasible to infer the missing information backwards based on a learned pattern if we can capture the shared high-level semantic information [19]. In light of above analysis, we design a cross-view imputation module with an autoencoder architecture where the encoder aims to learn high-level semantic representation across multi-view, while the decoder employs these cross-view fused high-level representation to reconstruct missing view information, guiding the model to fill data gaps.
For another, the given raw data usually come with many redundant features [34], and blindly using raw data directly to impute missing values may accumulate data redundancy. To address the challenges, drawing inspiration from the information bottleneck theory [35][36], we propose a task-oriented feature augmentation network to identify and select a subset of features that are most pertinent to the task, thereby ensuring that clean features are obtained before imputation.
@@ III-A Motivation
III-B Task-Oriented Feature Augmentation
When we tackle the complex problem of recovering missing multi-view data, we recognize that original data often contain task-irrelevant redundant information which may undermine the validity of the subsequent imputation process and the final classification results [3][32][35]. To this end, we seek to obtain high-quality discriminative and compact representation $V_{l}$ that densely encapsulates task-relevant information to replace the initial view redundant features $X_{l}$ before imputation. This target can be realized by minimizing the following problem according to the information bottleneck principle [37]:
@@ III-B Task-Oriented Feature Augmentation
$\frac{1}{nc}\sum_{i=1}^{n}\sum_{j=1}^{c}\left[\left(1-Y_{i,j}\right)\log{\left(1-Y^{{}^{\prime}}_{l_{i,j}}\right)}+Y_{i,j}\log{Y^{{}^{\prime}}_{l_{i,j}}}\right]$ (2)
@@
Then, for the second term $I\left({V_{l}};{X_{l}}\right)$, we introduce the variational encoding-decoding technique [40][41]. And to clarify the subsequent discussion more clearly, we’d better to start by explaining the basic concept of mutual information. That is, given ${V_{l}}$ and ${X_{l}}$, their joint probability distribution is denoted as $p(V_{l},X_{l})$ and their marginal probability distribution are denoted as $p(V_{l})$ and $p(X_{l})$, respectively, then the mutual information $I(V_{l};X_{l})$ is defined as:
$I\left(V_{l};X_{l}\right)=H\left(V_{l}\right)-H\left(V_{l}|X_{l}\right)$ (3)
$=\mathbb{E}_{p\left(V_{l},X_{l}\right)}\left[\log{\left(\frac{p\left(V_{l},X_{l}\right)}{p\left(V_{l}\right)p\left(X_{l}\right)}\right)}\right]$
@@
$\frac{1}{2nd_{l_{v}}}\sum_{i=1}^{n}\sum_{j=1}^{d_{l_{v}}}\left[(\mathcal{Q}_{l_{i,j}})^{2}-\log{(\mathcal{Q}_{l_{i,j}})^{2}}+(\mathcal{P}_{l_{i,j}})^{2}-1\right]$ (5)
@@
$L_{ta}^{l}$ $=\frac{1}{nc}\sum_{i=1}^{n}\sum_{j=1}^{c}\left[\left(1-Y_{i,j}\right)\log{\left(1-Y_{l_{i,j}}^{{}^{\prime}}\right)}\right.$ (6)
$\left.+Y_{i,j}\log{Y_{l_{i,j}}^{{}^{\prime}}}\right]U_{i,l}G_{i,j}+\frac{\delta}{2nd_{l_{v}}}\sum_{i=1}^{n}\sum_{j=1}^{d_{l_{v}}}$
$\left[(\mathcal{Q}_{l_{i,j}})^{2}-\log{(\mathcal{Q}_{l_{i,j}})^{2}}+(\mathcal{P}_{l_{i,j}})^{2}-1\right]U_{i,l}$
@@
III-C Cross-View Imputation

### v2

@@ Abstract
task-augmented, cross-view imputation, partial multi-view learning, incomplete multi-label classification
journal: Neural Networks
[…]
[address1]organization=College of Big Data and Information Engineering, Guizhou University,city=Guiyang, country=China
[…]
[address2]organization=Shenzhen Key Laboratory of Visual Object Detection and Recognition, Harbin Institute of Technology,city=Shenzhen, country=China
[…]
[address3]organization=School of Fashion and Textiles, The Hong Kong Polytechnic University,city=Hong Kong, country=China
[…]
[address4]organization=Hong Kong Laboratory for Artificial Intelligence in Design,city=Hong Kong, country=China
1 Introduction
@@
In recent years, with the significant progress of representation learning methods, traditional single-view data analysis methods are gradually showing their limitations, and they are difficult to fully adapt to the growing demands of diverse application scenarios [1][2][3]. Therefore, multi-view data integration has become an important way to break through the inherent barriers of single data sources and alleviate the lack of information [4][5]. This trend is accompanied by the rapid development of data acquisition technology, which makes a huge amount of multi-view data from all kinds of media or with different styles emerge, and providing more possibilities for achieving a comprehensive and accurate description of observation targets [6][7]. In short, multi-view data refers to a series of complementary information collections collected from different observation angles for the same object [8][9]. For example, three unique visual information obtained from X-rays, CT scans, and magnetic resonance imaging (MRI) in medical diagnostics constitute a multi-view dataset; retinal images acquired at four different positions can also constitute another multi-view retinal dataset [10]. More commonly, multi-view or multi-modal datasets that incorporate multiple types of data including text, images, and videos have been utilized in many application scenarios such as web page classification [11][12]. Concurrently, it is observed that single-label data frequently fail to adequately reflect the intricacies of real-world scenarios [13]. For instance, a picture of a flower may also require labels such as “winter” or “greenhouse”, demonstrating that multi-label provide a more comprehensive category feature space and more accurately preserve inter-label relationships [14]. Given the diversity of objects in multi-label classification tasks, researchers have innovatively combined multi-view learning with multi-label classification. This integration not only allows for the full utilization of data from various perspectives, but also provides the chance to capture the complex relationships within target objects and among labels [15]. These advantages collectively improve classification performance, making multi-view multi-label classification (MVMLC) a domain of significant research interest [16].
@@
For MVMLC, many methods have been proposed in the past few years, such as the latent semantic aware MvMLC [17], multi-view probabilistic model LCBM [18], and maximizing label-feature dependency based method [19]. However, MVMLC assumes all views and their corresponding labels are complete, a premise that does not always correspond to real-world situations [20][21]. For instance, in the context of video material, audio tracks are often absent, and text descriptions may be incomplete. Similarly, multimedia content disseminated on social media platforms frequently exhibits incomplete annotations owing to varying levels of user engagement. These practical situations give rise to the common challenge of incomplete multi-view features and incomplete category information in real-world datasets, which we define in our paper as “partial multi-view incomplete multi-label classification (PMVIMLC)” [22]. To solve such problem, various approaches for PMVIMLC have been proposed recent years. Tan et al. initially introduced the iMvWL model to construct a shared low-dimensional subspace based on matrix decomposition [16]. And Li et al. proposed the NAIM3L model to preserve the global consistency and local properties of the label structure [23]. Nevertheless, since these models employ shallow machine learning techniques to explore the information contained in features and labels, they are constrained in their ability to extract higher-order discriminative features from data. Therefore, Liu et al. developed the DICNet model for extracting deeper semantic representations from multi-view data [12]. However, the aforementioned models introduce a missing view indication matrix to circumvent unknown views and rely solely on the available data for classification, which fails to fully utilize the potential cross-view complementary information and the intrinsic relationships of the missing data, rendering it suboptimal [24]. To address this limitation, we propose to construct models that are capable of directly recovering missing data, despite the inherent challenges associated with this task.
@@
In light of such an issue, we propose a task-augmented cross-view imputation network (TACVI-Net) for partial multi-view incomplete multi-label classification. TACVI-Net has been designed to handle arbitrary view-missing scenarios and enhance representation learning through effective view completion. It is acknowledged that raw data may contain task-irrelevant redundant information [25], which may result in the accumulation of such information if used directly to fill in missing data. Therefore, we design a view-specific encoder-classifier architecture based on variational coding, inspired by the information bottleneck theory. This architecture can obtain enhanced features that are highly relevant to the task so that effectively avoiding the impact of irrelevant redundant information on model performance. In the absence of view features, it is of the utmost importance to determine an appropriate filling strategy to provide sufficient and useful information. Given the existence of the shared semantic information across multi-view, we convert each input view into a hidden representation via the view encoder and merge them into a joint representation to form a fused shared representation. Subsequently, the fused shared representation is employed to reconstruct missing data via the view decoder, which is then integrated with the existing data to construct a complete dataset. In this way, we can obtain more sufficient information to exploit the consistency and complementarity across different views, thus improving the final classification performance.
@@
2 Preliminaries
Table 1: Main notations utilized in our paper.
@@ 2 Preliminaries
$Y^{{}^{\prime}}$ predictive label matrix $(\left[0,1\right]^{n\times c})$
@@
$\mathbb{C}$ classifier for the second stage of fusion feature $\widehat{Z_{T}}$
2.1 Problem Statement and Associated Symbols
@@
For the PMVIMLC task, we delineate the following notations: $\left\{X_{l}\right\}_{l=1}^{m}$ and $Y$ symbolize the original multi-view data and corresponding labels, respectively, with $m$ denoting the total number of views. Each $X_{l}=\left\{x_{l_{1}},x_{l_{2}},...,x_{l_{n}}\right\}\in R^{n\times d_{l_{x}}}$ encapsulates the properties of instances within the $l$-th view, where $n$ represents the instance count and $d_{l_{x}}$ specifies the feature dimensionality for this view. Specifically, $x_{l_{i}}$ denotes the attributes of the $i$-th sample observed through the $l$-th perspective. The label matrix $Y=\left\{y_{1},y_{2},...,y_{n}\right\}\in\left\{0,1\right\}^{n\times c}$, with $c$ represents the label count, and $Y_{i,j}=1$ implies the $i$-th instance carries the $j$-th tag; conversely, $Y_{i,j}=0$. Specifically, $y_{i}$ represents the status of all possible labels for the $i$-th sample. To further account for missing data, we introduce a missing-view indication matrix $U\in\left\{0,1\right\}^{n\times m}$, where $U_{i,j}=1$ affirms the presence of the $i$-th sample in the $j$-th view, otherwise $U_{i,j}=0$. Similarly, $G\in\left\{0,1\right\}^{n\times c}$ serves as the missing-label indication matrix, with $G_{i,j}=1$ signifying awareness of the $j$-th label for the $i$-th instance, or $G_{i,j}=0$ for unknown status. In preprocessing, missing values in feature matrices $\left\{X_{l}\right\}_{l=1}^{m}$ and label matrix $Y$ are set as 0. Given the above complete or incomplete multi-view training data with complete or incomplete labels, the objective of our PMVIMLC task is to train a model that can accurately predict complete label sets for each multi-view sample with complete or incomplete views. To facilitate readers’ rapid identification of essential notations, we present a comprehensive list of the primary notations used in this paper in Table 1.
2.2 Related Works
2.2.1 Multi-view multi-label classification
@@
The field of MVMLC focuses on coping with data that contain multiple views and multiple labels at the same time which significantly increasing the complexity. Incorporating multi-view learning strategies has been shown to be more effective in solving such problems compared to multi-label classification approaches that rely only on a single view, inspiring a large increase of related research. For instance, Zhang et al. proposed a method that aligns the latent patterns of different views in the kernel space using matrix factorization, aiming to exploit the complementary information among multiple views [17]. Wu et al. introduced a method called SIMM, which employs orthogonal constraints on the shared subspace to extract unique information from multiple views, simultaneously optimizing the adversarial loss and the multi-label loss to obtain shared information across views [26]. Zhang et al. introduced RAIN, which facilitates interaction and semantic label embedding under the mechanism of multiple heads of attention to enhance label associations and dynamically acquire specific label representations based on learned label embedding [27]. Sun et al. put forward the generative model LCBM to fuse multi-view features and establish inter-label dependencies by using a Bernoulli mixture model [18].
@@
The CDMM method proposed by Zhao et al. employs a nonlinear kernel neural network to create separate classifiers for each view, ensuring inter-view consistency, and incorporates the Hilbert-Schmidt criterion to enhance inter-view diversity [28]. Yu et al. presented the M3AL method, which combines multi-view self-representation learning to optimize information separation, and employs a unique query strategy to select highly informative sample-label pairs for active learning based on shared, individual information and view differences [29]. Zhao et al. introduced the LVSL method for non-aligned multi-view multi-label classification, aiming to capture view-specific label information and low-level label structure under a unified system. This method exploits structural consistency between views and labels and enhances label spatial information [30]. The D-VSM approach proposed by Lyu et al. employs DeepGCN to directly encode feature representations at the individual view level, preserving the unique information of each view while facilitating effective coupling of information across views [31]. Yin et al. put forward the MMDOM method, which ensures that the low-dimensional shared subspace maintains both local and global data structures through first-order and second-order similarity matrices, effectively protecting manifold structures [32]. Xiao et al. developed an SVM-based approach, MVMLP, to obtain consensus information by reinforcing output similarity between views while introducing complementarity by using different views as privileged information for each other [33].
2.2.2 Partial multi-view incomplete multi-label classification
@@
Liu et al. proposed DICNet, which incorporates contrastive learning into the multi-view learning framework. This approach reduces the distance between positive sample pairs while increasing the distance between negative sample pairs, thereby enhancing the model’s capacity for information extraction across diverse views [12]. Liu et al. also proposed the MTD model, which decouples the common single-channel view-level representation in multi-view deep learning methods into shared and view-specific representations, with the aim of fully exploiting the complementary and consistent information across different views [37]. Tan et al. advocated for the application of information theory as a means to enhance and refine datasets that are inherently incomplete. The theoretical framework adopted by the researchers allows them to balance the available information to emphasize what is crucial for the task [38]. The SIP model, proposed by Liu and Xu et al., attempts to compress cross-view representations to maximize the amount of shared information, thereby better predicting semantic labels [39]. However, despite the diverse strategies offered by the aforementioned methods for addressing PMVIMLC, three key limitations persist: 1) High computational costs in training certain models. 2) Insufficient modeling of view-label interactions in some approaches, failing to capture complex dependencies. 3) Avoiding or eliminating missing view data, potentially leading to suboptimal information utilization. To overcome these issues, we propose to recover the information related to missing data by jointly exploring the interactions between multiple views and labels, and then using the recovered information as incremental information to break the performance limitations of existing methods.
Refer to captionRefer to caption
Figure 1: The sketch of our TACVI-Net. The TAEncoder refers to a task-augmented encoder based on variational encoding, while the VSEncoder and VSDecoder denote view-specific encoders and decoders that are designed with a symmetrical structure relative to each other, both of which are based on MLPs.
3 The proposed method
@@
3.1 Motivation
For one thing, it is well known that different views share the same high-level semantic information in the classification task because they are different descriptions of the same abstract target. This means that it becomes feasible to infer the missing information backwards based on a learned pattern if we can capture the shared high-level semantic information [24]. In light of above analysis, we design a cross-view imputation module with an autoencoder architecture where the encoder aims to learn high-level semantic representation across multi-view, while the decoder employs these cross-view fused high-level representation to reconstruct missing view information, guiding the model to fill data gaps. For another, the given raw data usually come with many redundant features [40], and blindly using raw data directly to impute missing values may accumulate data redundancy. To address the challenges, drawing inspiration from the information bottleneck theory [41][42], we propose a task-oriented feature augmentation network to identify and select a subset of features that are most pertinent to the task, thereby ensuring that clean features are obtained before imputation.
@@ 3.1 Motivation
3.2 Task-Oriented Feature Augmentation
When we tackle the complex problem of recovering missing multi-view data, we recognize that original data often contain task-irrelevant redundant information which may undermine the validity of the subsequent imputation process and the final classification results [38][41]. To this end, we seek to obtain high-quality discriminative and compact representation $V_{l}$ that densely encapsulates task-relevant information to replace the initial view redundant features $X_{l}$ before imputation. This target can be realized by minimizing the following problem according to the information bottleneck principle [43]:
@@ 3.2 Task-Oriented Feature Augmentation
$-I\left({V_{l}};Y\right)=\frac{1}{nc}\sum_{i=1}^{n}\sum_{j=1}^{c}\left[\left(1-Y_{i,j}\right)\log{\left(1-Y^{{}^{\prime}}_{l_{i,j}}\right)}+Y_{i,j}\log{Y^{{}^{\prime}}_{l_{i,j}}}\right]$ (2)
@@
Then, for the second term $I\left({V_{l}};{X_{l}}\right)$, we introduce the variational encoding-decoding technique [46]. And to clarify the subsequent discussion more clearly, we’d better to start by explaining the basic concept of mutual information. That is, given ${V_{l}}$ and ${X_{l}}$, their joint probability distribution is denoted as $p(V_{l},X_{l})$ and their marginal probability distribution are denoted as $p(V_{l})$ and $p(X_{l})$, respectively, then the mutual information $I(V_{l};X_{l})$ is defined as:
$I\left(V_{l};X_{l}\right)=H\left(V_{l}\right)-H\left(V_{l}|X_{l}\right)=\mathbb{E}_{p\left(V_{l},X_{l}\right)}\left[\log{\left(\frac{p\left(V_{l},X_{l}\right)}{p\left(V_{l}\right)p\left(X_{l}\right)}\right)}\right]$ (3)
@@
$I\left({V_{l}};{X_{l}}\right)=\frac{1}{2nd_{l_{v}}}\sum_{i=1}^{n}\sum_{j=1}^{d_{l_{v}}}\left[(\mathcal{Q}_{l_{i,j}})^{2}-\log{(\mathcal{Q}_{l_{i,j}})^{2}}+(\mathcal{P}_{l_{i,j}})^{2}-1\right]$ (5)
@@
$L_{ta}^{l}$ $=\frac{1}{nc}\sum_{i=1}^{n}\sum_{j=1}^{c}\left[\left(1-Y_{i,j}\right)\log{\left(1-Y_{l_{i,j}}^{{}^{\prime}}\right)}+Y_{i,j}\log{Y_{l_{i,j}}^{{}^{\prime}}}\right]U_{i,l}G_{i,j}$ (6)
$+\frac{\delta}{2nd_{l_{v}}}\sum_{i=1}^{n}\sum_{j=1}^{d_{l_{v}}}\left[(\mathcal{Q}_{l_{i,j}})^{2}-\log{(\mathcal{Q}_{l_{i,j}})^{2}}+(\mathcal{P}_{l_{i,j}})^{2}-1\right]U_{i,l}$
@@
3.3 Cross-View Imputation